OpenStudy (anonymous):

A net force of 100 N is moving with an acceleration of 5m/s2. What is the mass of the object? A.20 kg B.95 kg C.500 kg

OpenStudy (anonymous):

Fn=ma Fn-Net Force M- Mass A- acceleration In this case you fnet (net force) is 100 N , your acceleration is 5m/s2 and you mass is unknown. You can alter the formula into Fn/A=M which will give you 100/5= 20 Therefore your answer is 20 kg which is A

OpenStudy (anonymous):

Force = mass X acceleration Mass = Force / Acceleration F = 100 N A = 5m/s^2 M = 100/5